OpenStudy (anonymous):

My test has five question. But it looks very differnt from what we just did :(

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

ha

OpenStudy (anonymous):

FLVS?

OpenStudy (anonymous):

Identify whether the series \[\sum_{i=1}^{\infty} \] is a convergent or divergent geometric series and find the sum, if possible.

OpenStudy (anonymous):

and yes flvs

OpenStudy (anonymous):

yep. they always seem to do that. The tests have questions that u never see in the lesson. good luck Lesha.

ganeshie8 (ganeshie8):

@Jhannybean

OpenStudy (anonymous):

@kjuchiha ya ikr ! i just did the lesson and im like wow this is easy . and now the test im lik woaah this looks totally different then the lesson . and thankyouu (:

OpenStudy (jhannybean):

Where is your an?? I just see the starting of the series..

OpenStudy (anonymous):

what kind of series is that? dont have nth term

OpenStudy (anonymous):

oops soorryy. Identify whether the series \[\sum_{i=1}^{\infty}16(5)^{i-1}\] is a convergent or divergent geometric series and find the sum, if possible.

OpenStudy (jhannybean):

Oh okay. From the geometric series \[|r| < 1 \] for the series to be convergent.is r< 1?

OpenStudy (jhannybean):

\[\large \sum_{i=1}^{\infty}a_{1}r^{i-1}\]

OpenStudy (jhannybean):

\[\large \sum_{i=1}^{\infty}(16)(5)^{i-1}\]

OpenStudy (jhannybean):

Match up the terms to the definition of geometric series to help you out.

OpenStudy (anonymous):

ok so what does convergent and divergent mean ?

OpenStudy (jhannybean):

convergent means the behavior of the function goes to 0 as the function gets infinitely large and diverges means as the function approaches an infinite limit, the function will take on a very large value.

OpenStudy (anonymous):

okk . so how would i solve this problem ?

OpenStudy (jhannybean):

Observe that the formula for a geometric series \[\large \sum_{i=1}^{\infty}a_{1}r^{i-1}\]is where \(\large a_{1}\) is your first term and \(\large r\) is your ratio. Your equation is \[\large \sum_{i=1}^{\infty}16(5)^{i-1}\] where \(\large a_{1}= 16\) and \(\large r=5\). In order for this seriesto be convergent, the \(\large |r|\) has to be \(\large < 1\). in this case \(\large r = 5\) and therefore is not less than 1. Therefore the series is divergent. http://www.purplemath.com/modules/series5.htm